The ratio of men to women working for a company 2 to 3 . If there are 54 men working for the company, what is the total number of employees?

Answer by checkley79(3341)

2/3=54/X CROSS MULTIPLY.
2X=54*3
2X=162
X=162/2
X=81 WOMEN.
54+81=135 TOTAL EMPLOYEES.
PROOF:
2/3=54/81
3*54=2*81
162=162